What is the equation of a line which has 4 as x-intercept and -6 as y-intercept?

Grade:10

2 Answers

Mayank
6 years ago
As we know that intercept form of a line is given by the equation x/a+y/b=1..............(1)and you have given the x intercept as 4 and y intercept as -6so, put a=4 and b= -6 in equation 1x/4+y/-6=1(3x-2y)/12=13x-2y=12this is the correct answer.
sahil
6 years ago
the answer to this question is x/4-y/6=1 solving 3x-2y=12in y=mx+c form y=3/2x-6give approval if correct